OpenStudy (anonymous):

a line intercepts at the y axis at -7. if one point on the line the line is (4,5), what is the slope of the line

OpenStudy (calculusxy):

Use rise/run U have to connect the point on the y axis to the x axis. So to this first see how we can go vertically and then horizontally. U have to go up 11, making it to a positive. Then you have to to go to the right, 12. So it is rise/run=11/12 Now divide 11/12=0.916666....

OpenStudy (anonymous):

But is that the slope of the line?

OpenStudy (calculusxy):

Yes, I believe?

OpenStudy (anonymous):

i have a list of answer choices and they are: -1/3 -3 1/3 3

OpenStudy (calculusxy):

Sorry but that's so far learned

OpenStudy (anonymous):

@hartnn can you help me with this please

hartnn (hartnn):

y=mx+c x=4, y=5, c=-7 fond m

hartnn (hartnn):

find** m from 5=4m-7 can u ?

hartnn (hartnn):

@student901210

OpenStudy (anonymous):

u just add 7 to 5 and get 12 then divide by four. m=3

hartnn (hartnn):

correct! :) the slope is indded =3

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^